two cars, 325 kilometers apart, travel toward each other. one car travels at a rate of 70 km/h and the other at 60km/h. after how many hours will they meet?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

two cars, 325 kilometers apart, travel toward each other. one car travels at a rate of 70 km/h and the other at 60km/h. after how many hours will they meet?
---------------
The closing speed is (70+60)= 130 km/h
------------------
time = distance/rate
time = 325/130 = 2.5 hours.
==================================
Cheers,
Stan H.

Answer by richwmiller(17219)

70t+60t=325
rt=d
in this case
rt+rt=d
t=5/2=2.5 hours
check
2.5*70+2.5*60=325
140+35+120+30=325
260+65=325
325=325
ok